Accelerator Update for March 25, 2009 - March 27, 2009 |
Wednesday March 25 The day shift began with Operations monitoring the Tevatron (TeV) store 6927 (139.98E30) and the Recycler stash (141.4E10), with the Antiproton Source stacking (25.6E10), with Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. At 8:42 AM, a Booster RF station (BRF5) tripped off. It reset. At 8:45 AM, a Booster vacuum pump (IP61) tripped off for the first of three times in four hours. Operations prepared MI-31 for a 30-minute access by Pelletron physicists at 10 AM. At 10:03 AM, BRF16 tripped off. It reset. At 10:56 AM, the Recycler sequencer failed on the ACL script “RRSHOT.” There was an instruction to reboot a computer (DUE33) in the script. Operators rebooted it. A controls system administrator replaced DUE33 at 1:30 PM. A MI physicist began a study at 1:35 PM. Recycler physicists began a “mining” study at 2:03 PM. This study delayed antiproton transfers. At 3:16 PM, Operations reported more problems with DUE33. A controls person investigated. At 3:59 PM, the Booster beam permit tripped off with no indication of the cause. The Proton Source specialist, a controls person, and operators investigated. Operations reported at 4:02 PM that the Recycler had stashed antiprotons six times during the day shift. At 4:25 PM, BRF5 tripped off. It reset. A MI physicist completed his MI-8 study at 6:03 PM. Operations terminated store 6927 at 6:28 PM. Operations put the TeV into shot setup at 6:40 PM. At 7:28 PM, a Booster vacuum pump (IP61) tripped off. It reset. Operations established store 6928 around 8:26 PM with an initial luminosity of 275.8E30. Operations reported at 11:59 PM that the Recycler had stashed six times during the evening shift. Thursday March 26 The midnight shift began with Operations monitoring store 6928 (155.64E30) and the stash (128.7E10), with Pbar stacking (33.1E10), and with NuMI and MiniBooNE taking beam. At 12:32 AM, BRF7 and a Booster vacuum pump (IP122) tripped off. They reset. At 1:12 AM, Booster vacuum pumps IP61, 71, and 72 tripped off. They reset. At 4 AM, the P1 transfer line beam permit tripped off due to a power supply’s (I:V714) current drifting low. Everything reset. FESS personnel began work at 4:04 AM on the Booster East and West Gallery chilled water systems. At 7:45 AM, the P1 transfer line beam permit tripped off due to I:V714. Everything reset. At 7:49 AM, B:IP62 tripped off, followed by IP71 and 72. Everything reset. Operations terminated store 6928 at 8:04 AM. Operations reported at 8:09 AM that the Recycler had stashed eight times during the midnight shift. Cryo system technicians began an hour of work at 8:12 AM on the wet engines at TeV sectors B1 and A2. TeV power supply technicians began a 30-minute job at 8:13 AM on a TeV power supply (T:QAF4). Operations put the TeV through a dry squeeze at 9:20 AM. Operations put the TeV through a wet squeeze at 10:02 AM. Operations put the TeV into shot setup at 10:40 AM. Operations established store 6931 at 12:28 PM with an initial luminosity of 300.9E30. At 1:13 PM, Operations disabled the MiniBooNE autotune program. It was not setting devices correctly. At 3:18 PM, the P1 transfer line beam permit tripped off when I:V714’s current drifted low. Everything reset. Operations reported at 3:58 PM that the Recycler had stashed five times during the day shift. At 4:57 PM, three Booster vacuum pumps (IP61, 62, and 72) tripped off. They all reset. At 6:15 PM, I:V714 tripped off. It reset. At 9:12 PM, LRF3 tripped off. It reset. At 9:45 PM, operators reset the 186 card for MIVAC. It had stopped reporting. FESS personnel completed their chilled water work in the Booster West Gallery at 10:18 PM. At 11:03 PM, Operations lowered the intensity for NuMI due to bad beam positions in MI and Booster. Operations reported at midnight that the Recycler had stashed nine times during the evening shift. Friday March 27 The midnight shift began with Operations monitoring store 6931 (90.68E30) and the stash (317.3E10), with Pbar stacking (32.2E10), and with NuMI and MiniBooNE taking beam. Operations terminated store 6931 at 1:47 AM. Operations put the TeV into shot setup at 2:28 AM. At 2:32 AM, BRF2 tripped off for the first of three times in three minutes. It reset. Operations established store 6932 at 3:31 AM with an initial luminosity of 306.76E30. The Neutron Therapy Facility reported at 7:43 AM that they had begun treating the first of three patients for the day. At 7:55 AM, Operations reported that they rebooted a Pbar Accumulator computer because of stale data for ARF1. It did not help. Further investigation found a problem with the VSA. It was restarted. Operations reported at 8:05 AM that the Recycler had stashed six times during the midnight shift. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/27/2009 email Fermilab |
| Security, Privacy, Legal | |